Write the expression as a complex number in standard form.
(2+i)+(3+2i)

Found 2 solutions by stanbon, gonzo:
Answer by stanbon(48545)

Write the expression as a complex number in standard form.
(2+i)+(3+2i)
= 5 + 3i
=============
Cheers,
Stan H.

Answer by gonzo(654)

you add the real parts together and you add the imaginary parts together.
2 + 3 = 5
i + 2i = 3i
answer is:
5 + 3i
-----